OpenStudy (anonymous):

PLEASE help me solve for g !!! :) -10 = -2 + 1/2 d

OpenStudy (anonymous):

add 2 to both sides: -8=1/2d multiply by 2: -16=d (or g whatever variable you want)

OpenStudy (anonymous):

d=5

OpenStudy (anonymous):

@ Jamesumner93 : can you help me with another problem it is a little bit different and i am just curious about if i solve it the same way

OpenStudy (anonymous):

5x - 11 = 7x - 3 help me solve for x

OpenStudy (anonymous):

add 11 to both sides: 5x=7x+8 then subtract 7x (trying to get x's on one side): -2x=8 then divide by -2: x=-4